How do you solve for APR in this equation?
1.0593 = (1 + APR/3.3182)) ^3.3182

Answer by Theo(13342)

looks like you caught a tiger by the tail.

the easiest way to solve this is to let x = apr and let y = (1 + (x/3.3182))

your equation becomes:

1.0593 = y^3.3182

you can take the 3.3182 root of both sides of this equation to get:

1.0593^(1/3.3182) = y

use your calculator to get y = 1.017512898

your equation of 1.0593 = y^3.3182 is now equal to:

1.0593 = 1.017512898^(3.3182)

we original set y = 1 + x/3.3182

we know that y = 1.017512898

we can substitute for y to get:

1.017512898 = 1 + (x/3.3182)

subtract 1 from both sides of this equation to get:

.017512898 = x/3.3182

multiply both sides of this equation by 3.3182 to get:

x = .058111297 = apr

that should be your answer.

to confirm, plug that into your original equation to see if it is true.

your original equation is:

1.0593 = (1 + (APR/3.3182)) ^3.3182

replace apr with .058111297 to get:

1.0593 = (1 + (.058111297/3.3182)) ^(3.3182)

simplify to get:

1.0593 = 1.017512898^(3.3182)

simplify further to get:

1.0593 = 1.0593 confirming that the value of your apr that we calculated is good.

in your equation of:

1.0593 = y^3.3182

you could also have taken the log of both sides of the equation to get:

log(1.0593) = log(y^3.3182) which becomes:

log(1.0593) = 3.3182 * log(y)

divide both sides of the equation by 3.3182 to get:

log(1.0593) / 3.3182) = log(y)

simplify by taking the log of 1.0593 and dividing by 3.3182 to get:

log(y) = .007539923

take the anti-log of both sides of the equation to get:

y = 1.017512898

either way would have gotten you to y = 1.017512898